The marginal propensity to consume (MPC) is 0.60 The multiplier is 25 (Round your answer to one decimal place) Suppose that net exports changes by $-100 The change in real GDP will be S(Round your answer to the nearest dollar)
Answer #1

Answer'Multilier=1/(1-MPC)
=1/(1-0.6)
=2.5
the multiplier is 2.5
the change in GDP because of exports=multiplier*change in exports
=2.5*(-100)
=$-250
The Change in real GDP will be $-250
means the GDP decreases by $250